Fortes Inc. has provided the following data concerning one of the products in its standard cost system. Variable manufacturing overhead is applied to products on the basis of direct labor-hours.

Inputs Standard Quantity or Hours per Unit of Output Standard Price or Rate
Direct materials 7.1 ounces $ 7.30 per ounce
Direct labor 0.6 hours $ 26.70 per hour
Variable manufacturing overhead 0.6 hours $ 5.60 per hour

The company has reported the following actual results for the product for April:

Actual output 8,550 units
Raw materials purchased 61,710 ounces
Actual cost of raw materials purchased $ 341,700
Raw materials used in production 60,735 ounces
Actual direct labor-hours 3,850 hours
Actual direct labor cost $ 107,860
Actual variable overhead cost $ 20,701

Required:

a. Compute the materials price variance for April.

b. Compute the materials quantity variance for April.

c. Compute the labor rate variance for April.

d. Compute the labor efficiency variance for April.

e. Compute the variable overhead rate variance for April.

f. Compute the variable overhead efficiency variance for April.

(Indicate the effect of each variance by selecting "F" for favorable, "U" for unfavorable, and "None" for no effect (i.e., zero variance). Input all amounts as positive values.)

a Materials price variance
b Materials quantity variance
c Labor rate variance
d Labor efficiency variance
e Variable overhead rate variance
f Variable overhead efficiency rate
Answer #1

Direct Material Price Variance = (Standard price – Actual Price)*Actual Quantity purchased

= (7.30 – Actual Rate)*61,710

= 7.30*61710 - 341700

= 108,783 Favorable

Direct Material Quantity Variance = (Standard Quantity – Actual Quantity)*Standard Price

= (8550*7.1 – 60735)*7.30

= $219 unfavorable

Direct labour rate variance = (Standard rate – actual rate)*actual hours

= (26.70-Actual rate)*3,850

= 26.7*3850 - 107860

= $5,065 unfavorable

Direct Labor Efficiency Variance = (Standard Hours – Actual Hours)*Standard rate

= (8550*0.6 – 3850)*26.70

= $34,176 favorable

Variable overhead rate = (5.60-Actual)*3850

= 5.60*3850 – 20701

= $859 favorable

Efficiency = (8550*0.6-3850)*5.60

= $7168 favorable